# Learning to Defer in Content Moderation: The Human-AI Interplay

### v2

@@ 1.1 Our contributions
  • First, existing content moderation practice typically uses static thresholds on the sample average of a post’s harmfulness to decide which posts to admit for human reviews (see Appendix A). We demonstrate that such approaches can be inefficient by failing to consider the variance of post harmfulness and the fluctuation in post arrivals and human capacity prominently seen in practice (see [MSA+21, Figure 2].) Instead, we propose a congestion-aware admission rule that balances posts’ idiosyncrasy loss when we rely on AI prediction and the additional congestion in the human review system caused by admitting a new post.
@@
To formally tackle the problem, we introduce a model that combines learning to defer with soft capacity constraints via queueing delays. To the best of our knowledge, the impact of limited capacity for learning to defer has not been considered in the literature (see Section 1.2 for further discussion). In particular, in each period $t=1,\ldots,T$, a post $j$ arrives with a type $k(j)$ drawn from a time-varying distribution over $K$ types. A post $j$ has an unknown cost $c_{j}=h_{k(j)}+\xi_{j}$ where $\xi_{j}\sim\mathcal{N}(0,\sigma_{k(j)}^{2})$ is a zero-mean idiosyncratic Gaussian noise while $h_{k}$ and $\sigma_{k}^{2}$ are the expectation and variance of the cost of type-$k$ posts. To capture the difference in reliability, we assume that the AI model only observes the type of the post; the true cost $c_{j}$ is only observable via human reviews. To capture the impact of limited human capacity, we assume that posts that are admitted (deferred) by the AI wait in a review queue. At the end of a period, one post in the queue is scheduled for review; the review is completed in this period with a probability $\mu_{k}N(t)$ that depends on both the specific post type $k$ and the fluctuating capacity of reviewers $N(t)$. We assume that a type-$k$ post is viewed only within $\ell_{k}$ periods since its arrival and only if it is not removed.
@@
As typical in complex learning settings, we start by assuming knowledge of the latent parameters (i.e., the expected costs $h_{k}$). The classification decision is then to keep the post $j$ if and only if the expected cost of its type is negative: $h_{k(j)}<0$. Of course, deferring a post to human reviewers can improve upon this ex-ante classification decision as human reviewers can observe the true cost. The admission decision should thus identify the posts that would benefit the most from this more refined observation. If all types have equal variance then the human reviewing capacity is more efficiently allocated when it focuses on borderline posts, i.e., $h_{k(j)}$ close to 00, as borderline types have higher probability that the true cost of their posts has a different sign from their expected cost (which would lead to a different classification by the omniscient benchmark). This intuition drives the design of admission policies in existing content moderation practice [MSA+21, ABB+22] that operate based on two thresholds: a post $j$ is rejected (resp. accepted) without admission if its expected cost $h_{k(j)}$ is above the higher (resp. below the lower) threshold and is admitted for human review if it lies between the two thresholds. A more detailed comparison to these works is provided in Appendix A.
@@
That said, even when all expected costs $h_{k}$ are known, there are two important shortcomings in these two-threshold heuristics. First, when the variance of different types is heterogeneous, the admission decision should not be restricted to the expected costs but rather also take variance into account. In particular, it may be more beneficial to defer to human reviewers post from a non-borderline type $k$ with high idiosyncratic variance $\sigma_{k}^{2}$ instead of wasting human capacity on more deterministic borderline types. Second, even if we operate with variance-aware scores, using static thresholds does not allow the system to adapt to the time-varying arrival patterns of posts and the fluctuating human capacities that arise in practice (see [MSA+21, Figure 2]). In particular, in certain time periods, the number of borderline posts may be either too large resulting in high delays for admitted posts due to the overflow in the review system or too small resulting in misclassification for some non-borderline posts that could have been prevented if they were admitted.
@@
Our approach (Section 3) quantifies and balances the two losses hinted above (due to idiosyncrasy and delay) and is henceforth called Balanced Admission Control for Idiosyncrasy and Delay or BACID as a shorthand. We first calculate the expected ex-ante per-period loss of keeping a post on the platform (resp. removing it) using the distributional information of the true cost $c_{j}$; this loss is denoted by $r_{k}^{O}$ (resp. $r_{k}^{R}$). Hence, if a post is not admitted to human review, its per-period loss is $r_{k}=\min(r_{k}^{O},r_{k}^{R})$ and thus its idiosyncrasy loss that aggregates over its lifetime is $r_{k}\ell_{k}$. On the other hand, admitting a post incurs delay loss, which is the increase in congestion for future posts due to the limited human capacity. Motivated by the drift-plus-penalty literature [Nee22], we estimate the delay loss of a post by the number of posts of the same type currently waiting for reviews. We then admit a post if (a weighted version of) its idiosyncrasy loss (which one can avoid by admission) exceeds the estimated delay loss (which occurs due to the admission). For scheduling, we choose in each period the type with the most number of waiting posts (weighted by the difficulty to review this type) to review. We show that while heuristics like AI-Only or Human-Only have average regret at least $\Omega(\ell_{\max})$ (Propositions 1 and 2), where $\ell_{\max}$ is the maximum lifetime, BACID achieves a $O(\sqrt{K\ell_{\max}})$ average regret even with general time-varying arrivals and capacities (Theorem 1). This bound is optimal on the dependence on $\ell_{\max}$ as we also show a lower bound $\Omega(\sqrt{\ell_{\max}})$ for any deterministic policy (Theorem 2).
@@
To deal with unknown expected costs $h_{k}$, a typical approach in the bandit literature is to use an optimistic estimator for the unknown quantities. The unknown quantity that affects our admission rule is $r_{k}$, which depends on the unknown expected cost $h_{k}$. Following the optimistic approach, we first create a confidence interval $[\underline{h}_{k},\bar{h}_{k}]$ for the unknown expected cost $h_{k}$ based on reviewed posts. We then compute an optimistic estimator $\bar{r}_{k}$ using the parameter $\widetilde{h}_{k}\in[\underline{h}_{k},\bar{h}_{k}]$ that maximizes $r_{k}(\widetilde{h}_{k})$.The optimistic admission rule then admits a new post if (a weighted version of) the optimistic idiosyncrasy loss, $\bar{r}_{k}\ell_{k}$, exceeds its estimated delay loss. The high-level idea behind optimism is that, if a post is admitted (due to the optimistic estimation), we collect additional samples for this post which result in shrinking the confidence interval and eventually leading the confidence interval to converge to the true parameter $h_{k}$. Without considering heterogeneity in variance, [ABB+22] adopts a similar optimism-only heuristic to address the online learning nature of $h_{k}$.
@@
However, optimism-only learning heuristics disregard the selective sampling nature of the feedback. When there is uncertainty on $h_{k}$, the benefit of admitting a post is not restricted to avoiding the idiosyncrasy loss of the current post, but extends to having one more data points that can improve the classification decision of future posts. This benefit does not arise when $h_{k}$ is known because the best classification decision is then clear (keep a post if and only if $h_{k}\leq 0$). With unknown $h_{k}$, there is an additional classification loss due to incorrect estimation of the sign of $h_{k}$. Given that optimism-only heuristics emulate the admission rule with known $h_{k}$, they disregard this classification loss and more broadly the positive externality that labels for the current post have on future posts. This challenge highlights that optimism-only heuristics such as Upper Confidence Bound (UCB) do not employ forward-looking exploration but rather optimistically myopic exploitation. They are myopic in their nature: they create a confidence interval around the parameter of interest and are optimistic with respect to this confidence interval, but they make an exploit decision based on those optimistic estimates. Concretely, given the most optimistic estimates about the underlying parameters, these algorithms select the action that myopically maximizes the contribution of the current post without actively considering any positive externality of the labels on future posts. We formally illustrate this inefficiency of optimism-only learning heuristics in Proposition 3.
@@
To circumvent the above issue, we design an online learning version of BACID (Section 4) , which we term OLBACID, that augments the optimistic admission with label-driven admission and forced scheduling. When a post arrives, we estimate its potential classification loss based on its current confidence interval; if this is higher than a threshold, we admit the post to a label-driven queue and prioritize posts in this queue via forced scheduling to ensure enough labels for future classification decisions. To avoid exhausting human capacity, we limit the number of posts in the label-driven queue to one at any time. We show that OLBACID enjoys an average regret of $\tilde{O}(\sqrt{K\ell_{\max}}+K\ell_{\max}/\sqrt{T})$ (Theorem 3), addressing the inefficiency of optimism-only heuristics.
@@
To remove the dependence on $K$, which can be large when types are defined based on contextual information (e.g., word embeddings of the contents), we provide a contextual bandit extension of our algorithm. In particular, we consider a linear contextual setting where $h_{k}=\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star}$ where $\boldsymbol{\phi}_{k}$ is a known $d-$dimensional feature vector for each type $k$ and $\boldsymbol{\theta}^{\star}$ is an unknown $d-$dimensional vector; this linear cost structure is similar to the one used in [ABB+22].
@@
Our instance-dependent guarantee $\tilde{O}(\sqrt{G\ell_{\max}}+\ell_{\max}(\Delta+d\sqrt{G/T})$ (Theorem 4) scales optimally with the maximum lifetime $\ell_{\max}$ and avoids the dependence on $K$. We also provide a worst-case guarantee $\tilde{O}(\ell_{\max}^{2/3}+\ell_{\max}^{7/6}d/\sqrt{T})$ with no dependence on $G$ and $\Delta$ at the cost of a worse dependence on $\ell_{\max}$ (Corollary 1). To the best of our knowledge, this is the first result for online learning in contextual queueing systems and hence our analytical framework may be of independent interest.
@@ 1.2 Related work
The nature of human-AI collaboration can be broadly classified into two types: augmentation and automation [BM17, RK21]. In particular, augmentation represents a “human-in-the-loop” type workforce where human experts combine machine learning with their own judgement to make better decisions. Such augmentation is often found in high-stake settings such as healthcare [LLL22], child maltreatment hotline screening [DFC20], refugee resettlement [BP22, AGP+23] and bail decisions [KLL+18]. A general concern for augmentation relates to humans’ compliance patterns to [BBS21, LLL22], the impact of such patterns to decision accuracy [KLL+18, MS22] and the impact on fairness [MSG22, GBB23]. Automation, on the other hand, concerns the use of machine learning in place of humans and is widely applied in human and social services where colossal demands overwhelm limited human capacity, such as in data labelling [VLSV23], content moderation [GBK20, MSA+21] and insurance [Eub18]. Full automation is clearly undesirable in these applications as machine learning can err. A natural question is how one can better utilize the limited human capacity when machine is uncertain about the prediction. The literature on “learning to defer”, which studies when machine learning algorithms should defer decisions to downstream experts, tries to answer this question and is where our work fits in.
@@ 2 Model
The per-view harmfulness of a post is captured by its cost $c_{j}\in\mathbb{R}$; this quantity being negative means that the post is healthy. Condition on having type $k$, the cost of post $j$ is given by $c_{j}=h_{k(j)}+\xi_{j}$ where $h_{k}$ is the average cost of type-$k$ posts and $\xi_{j}\sim\mathcal{N}(0,\sigma_{k}^{2})$ is independent Gaussian noise. Although the type $k(j)$ is observable, we assume the cost $c_{j}$ remains unknown until it is reviewed by a human. We next define the total harmfulness of a post. For every post of type $k$, we assume it has a lifetime $\ell_{k}$ and a per-period view $v_{k}$, i.e., if a post $j$ arrives in period $t(j)$, then for periods $\{t(j),\ldots,\min(t(j)+\ell_{k}-1,T)\}$, it will receive $v_{k}$ views given that it is on the platform. [^2] Our analysis easily extends to the case where the view per period is a random variable with mean $v_{k}$. We define $O_{j}(t)\in\{0,1\}$ such that it is equal to $1$ when post $j$ is on the platform for period $t$. The total harmfulness of a post $j$ is given by $c_{j}v_{k}\sum_{t=t(j)}^{\min(t(j)+\ell_{k}-1,T)}O_{j}(t)$. If $c_{j}$ is known, this quantity is minimized by setting $O_{j}(t)=1$ if $c_{j}\leq 0$ or 00 if $c_{j}>0$.

### v3

@@ 1.1 Our contributions
  • First, existing content moderation practice typically uses static thresholds on the sample average of a post’s harmfulness to decide which posts to admit for human reviews (see Appendix A). We demonstrate that such approaches can be inefficient by failing to consider the uncertainty of post harmfulness and the fluctuation in post arrivals and human capacity prominently seen in practice (see [MSA+21, Figure 2].) Instead, we propose a congestion-aware admission rule that balances posts’ idiosyncrasy loss when we rely on AI prediction and the additional congestion in the human review system caused by admitting a new post.
@@
To formally tackle the problem, we introduce a model that combines learning to defer with soft capacity constraints via queueing delays. To the best of our knowledge, the impact of limited capacity for learning to defer has not been considered in the literature (see Section 1.2 for further discussion). In particular, in each period $t=1,\ldots,T$, a post $j$ arrives with a type $k(j)$ drawn from a time-varying distribution over $K$ types. The type of a post can be viewed as extracted features about the post, e.g. the final layer of a neural network given the unstructured post as input or a word embedding indicating whether a the post contains a sensitive word.
To capture the difference in reliability between AI and humans, we employ the concept of aleatoric and epistemic uncertainty in machine learning [HW21]. In particular, each type $k$ has an unknown sub-Gaussian per-period cost distribution $\mathcal{F}_{k}$ with mean $h_{k}$ and a post $j$ has an unknown cost $c_{j}$ independently sampled from $\mathcal{F}_{k(j)}$. We assume that the AI model only observes the type of the post; the true cost $c_{j}$ is only observable via human reviews. This distinction captures the imperfect predictive power of extracted features (and thus the AI). On the one hand, having more data can help AI to learn the cost distributions $\mathcal{F}_{k}$, corresponding to epistemic uncertainty that is reducible by more data. On the other hand, even if the AI has assess to infinitely many data, it cannot exactly predict the cost of a post due to the randomness in $\mathcal{F}_{k}$ (aleatoric uncertainty conditioned on the extracted features). As a model assumption, humans can observe the true cost of a post because of having more cognitive abilities and not being limited by the extracted features.
To capture the impact of limited human capacity, we assume that posts that are admitted (deferred) by the AI wait in a review queue. At the end of a period, one post in the queue is scheduled for review; the review is completed in this period with a probability $\mu_{k}N(t)$ which depends on both the type-dependent per-reviewer service rate $\mu_{k}$  [^2] Service rates can be different across types. For example, it may be faster to review a text than a video post. and the fluctuating capacity of reviewers $N(t)$. We assume that a type-$k$ post is viewed only within $\ell_{k}$ periods since its arrival and only if it is not removed.
@@
As typical in complex learning settings, we start by assuming knowledge of the latent parameters (i.e., the cost distributions $\mathcal{F}_{k}$). The classification decision is then to keep the post $j$ if and only if the expected cost of its type is negative: $h_{k(j)}<0$. Of course, deferring a post to human reviewers can improve upon this ex-ante classification decision as human reviewers can observe the true cost. The admission decision should thus identify the posts that would benefit the most from this more refined observation. If all types have equal variance and costs are normally distributed, the human reviewing capacity is more efficiently allocated when it focuses on borderline posts, i.e., $h_{k(j)}$ close to 00, as borderline types have higher probability that the true cost of their posts has a different sign from their expected cost (which would lead to a different classification by the omniscient benchmark). This intuition drives the design of admission policies in existing content moderation practice [MSA+21, ABB+22] that operate based on two thresholds: a post $j$ is rejected (resp. accepted) without admission if its expected cost $h_{k(j)}$ is above the higher (resp. below the lower) threshold and is admitted for human review if it lies between the two thresholds. A more detailed comparison to these works is provided in Appendix A.
@@
That said, even when all expected costs $h_{k}$ are known, there are two important shortcomings in these two-threshold heuristics. First, when the cost distribution of different types is heterogeneous, the admission decision should not be restricted to the expected costs but rather also take the cost distribution into account. In particular, it may be more beneficial to defer to human reviewers a post from a non-borderline type $k$ with high idiosyncratic uncertainty instead of wasting human capacity on more deterministic borderline types. Second, even if we operate with distribution-aware scores, using static thresholds does not allow the system to adapt to the time-varying arrival patterns of posts and the fluctuating human capacities that arise in practice (see [MSA+21, Figure 2]). In particular, in certain time periods, the number of borderline posts may be either too large resulting in high delays for admitted posts due to the overflow in the review system or too small resulting in misclassification for some non-borderline posts that could have been prevented if they were admitted.
@@
Our approach (Section 3) quantifies and balances the two losses hinted above (due to idiosyncrasy and delay) and is henceforth called Balanced Admission Control for Idiosyncrasy and Delay or BACID as a shorthand. We first calculate the expected ex-ante per-period loss of keeping a post on the platform (resp. removing it) using the distributional information of the true cost $c_{j}$; this loss is given by $r_{k}^{O}\coloneqq\mathbb{E}\left[c_{j}^{+}\,\middle|\,k(j)=k\right]$ (resp. $r_{k}^{R}\coloneqq\mathbb{E}\left[-c_{j}^{-}\,\middle|\,k(j)=k\right]$). Hence, if a post is not admitted to human review, its per-period loss is $r_{k}=\min(r_{k}^{O},r_{k}^{R})$ and thus its idiosyncrasy loss that aggregates over its lifetime is $r_{k}\ell_{k}$. On the other hand, admitting a post incurs delay loss, which is the increase in congestion for future posts due to the limited human capacity. Motivated by the drift-plus-penalty literature [Sto05, HN11, Nee22], we estimate the delay loss of a post by the number of posts of the same type currently waiting for reviews. We then admit a post if (a weighted version of) its idiosyncrasy loss (which one can avoid by admission) exceeds the estimated delay loss (which occurs due to the admission). For scheduling, we choose in each period the type with the most number of waiting posts (weighted by the difficulty to review this type) to review. We show that while heuristics like AI-Only or Human-Only have average regret at least $\Omega(\ell_{\max})$ (Propositions 1 and 2), where $\ell_{\max}$ is the maximum lifetime, BACID achieves a $O(\sqrt{K\ell_{\max}})$ average regret even with general time-varying arrivals and capacities (Theorem 1). This bound is optimal on the dependence on $\ell_{\max}$ as we also show a lower bound $\Omega(\sqrt{\ell_{\max}})$ for any deterministic policy (Theorem 2). Numerical simulations demonstrate substantial improvement of BACID compared with congestion-unaware heuristics (Section 6.1), even when the latter heuristics can adapt to the non-stationary arrival and capacity patterns.
@@
To deal with unknown distributions, a typical approach in the bandit literature is to use an optimistic estimator for the unknown quantities. The unknown quantity that affects our admission rule is $r_{k}$, which depends on the unknown expected cost of keeping (resp. removing) a post $r_{k}^{O}$ (resp. $r_{k}^{R}$). Following the optimistic approach, we first create optimistic estimators $\bar{r}_{k}^{O},\bar{r}_{k}^{R}$ for the unknown expected costs $r_{k}^{O},r_{k}^{R}$ based on reviewed posts. We then compute an optimistic estimator $\bar{r}_{k}=\min(\bar{r}_{k}^{O},\bar{r}_{k}^{R})$. The optimistic admission rule then admits a new post if (a weighted version of) the optimistic idiosyncrasy loss, $\bar{r}_{k}\ell_{k}$, exceeds its estimated delay loss. The high-level idea behind optimism is that, if a post is admitted (due to the optimistic estimation), we collect additional samples for this post which result in shrinking the confidence interval and eventually leading the optimistic estimators to converge to the true parameters. Without considering heterogeneity in cost distributions, [ABB+22] adopts a similar optimism-only heuristic to address the online learning nature of the mean cost $h_{k}$.
@@
However, optimism-only learning heuristics disregard the selective sampling nature of the feedback. When there is uncertainty on cost distributions, the benefit of admitting a post is not restricted to avoiding the idiosyncrasy loss of the current post, but extends to having one more data points that can improve the classification decision of future posts. This benefit does not arise when cost distributions are known because the best classification decision is then clear —keep a post if and only if the expected cost of keeping it is lower, i.e., $r_{k}^{O}<r_{k}^{R}$. With unknown distributions, there is an additional classification loss due to incorrect estimation of the sign of $r_{k}^{O}-r_{k}^{R}$. Given that optimism-only heuristics emulate the admission rule with known distributions, they disregard this classification loss and more broadly the positive externality that labels for the current post have on future posts. This challenge highlights that optimism-only heuristics such as Upper Confidence Bound (UCB) do not employ forward-looking exploration but rather optimistically myopic exploitation. They are myopic in their nature: they create a confidence interval around the parameter of interest and are optimistic with respect to this confidence interval, but they make an exploit decision based on those optimistic estimates. Concretely, given the most optimistic estimates about the underlying parameters, these algorithms select the action that myopically maximizes the contribution of the current post without actively considering any positive externality of the labels on future posts. We illustrate this inefficiency of optimism-only learning heuristics formally in Proposition 3 and numerically in Section 6.2.
@@
To circumvent the above issue, we design an online learning version of BACID (Section 4) , which we term OLBACID, that augments the optimistic admission with label-driven admission and forced scheduling. When a post arrives, we estimate its potential classification loss based on its current confidence interval; if this is higher than a threshold, we admit the post to a label-driven queue and prioritize posts in this queue via forced scheduling to ensure enough labels for future classification decisions. To avoid exhausting human capacity, we limit the number of posts in the label-driven queue to one at any time. We show that OLBACID enjoys an average regret of $\tilde{O}(\sqrt{K\ell_{\max}}+K\ell_{\max}/\sqrt{T})$(Theorem 3) and has better numerical performance (Section 6.2), addressing the inefficiency of the optimism-only heuristic.
@@
To remove the dependence on $K$, which can be large when types are defined based on contextual information (e.g., word embeddings of the contents), we provide a contextual bandit extension of our algorithm. In particular, we consider a linear contextual setting where $[r_{k}^{O},r_{k}^{R}]=\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\Theta}^{\star}$ where $\boldsymbol{\phi}_{k}$ is a known $d-$dimensional feature vector for each type $k$ and $\boldsymbol{\Theta}^{\star}$ is an unknown $d\times 2$ dimensional matrix.
@@
Our instance-dependent guarantee $\tilde{O}(\sqrt{G\ell_{\max}}+\ell_{\max}(\Delta+d\sqrt{G/T}))$ (Theorem 4) scales optimally with the maximum lifetime $\ell_{\max}$ and avoids the dependence on $K$. We also offer a worst-case guarantee $\tilde{O}(\ell_{\max}^{2/3}+\ell_{\max}^{7/6}d/\sqrt{T})$ with no dependence on $G$ and $\Delta$ at the cost of a worse dependence on $\ell_{\max}$ (Corollary 1). To the best of our knowledge, this is the first result for online learning in contextual queueing systems and hence our analytical framework may be of independent interest.
@@ 1.2 Related work
The nature of human-AI collaboration can be broadly classified into two types: augmentation and automation [BM17, RK21]. In particular, augmentation represents a “human-in-the-loop” type workforce where human experts combine machine learning with their own judgement to make better decisions. Such augmentation is often found in high-stake settings such as healthcare [LLL22], child maltreatment hotline screening [DFC20], refugee resettlement [BP22, AGP+23] and bail decisions [KLL+18]. A general concern for augmentation relates to humans’ compliance patterns [BBS21, LLL22], the impact of such patterns to decision accuracy [KLL+18, MS22] and the impact on fairness [MSG22, GBB23]. Automation, on the other hand, concerns the use of machine learning in place of humans and is widely applied in human and social services where colossal demands overwhelm limited human capacity, such as in data labelling [VLSV23], content moderation [GBK20, MSA+21] and insurance [Eub18]. Full automation is clearly undesirable in these applications as machine learning can err. A natural question is how one can better utilize the limited human capacity when machine is uncertain about the prediction. The literature on “learning to defer”, which studies when machine learning algorithms should defer decisions to downstream experts, tries to answer this question and is where our work fits in.
@@ 2 Model
The per-period harmfulness of a post is captured by its cost $c_{j}\in\mathbb{R}$; this quantity being negative means that the post is healthy. [^3] One can interpret this cost as number of views times the per-view cost for each period. We make a simplifying assumption that this product is unchanged throughout the lifetime of the post. Condition on having type $k$, the cost $c_{j}$ of post $j$ is independently sampled from an unknown sub-Gaussian distribution $\mathcal{F}_{k}$ with mean $h_{k}$. Although the type $k(j)$ is observable, we assume the cost $c_{j}$ remains unknown until it is reviewed by a human. We use $\mathbb{E}_{k}\left[\cdot\right]$ to condition on a post type $k$, i.e., $\mathbb{E}_{k}\left[\cdot\right]=\mathbb{E}\left[\cdot\,\middle|\,k(j)=k\right]$.
We next define the total harmfulness of a post. For every post of type $k$, we assume it has a lifetime $\ell_{k}$ and a per-period view $v_{k}$, i.e., if a post $j$ arrives in period $t(j)$, then for any period in $\{t(j),\ldots,\min(t(j)+\ell_{k}-1,T)\}$, it adds a cost $c_{j}$ given that it is on the platform. We define $O_{j}(t)\in\{0,1\}$ such that it is equal to $1$ when post $j$ is on the platform for period $t$. The total harmfulness of a post $j$ is given by $c_{j}\sum_{t=t(j)}^{\min(t(j)+\ell_{k}-1,T)}O_{j}(t)$. If $c_{j}$ is known, this quantity is minimized by setting $O_{j}(t)=1$ if $c_{j}\leq 0$ or 00 if $c_{j}>0$.